# CSRM-LLM: Embracing Multilingual LLMs for Cold-Start Relevance Matching in Emerging E-commerce Markets

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761545isbn: 979-8-4007-2040-6/2025/11ccs: Information systemsccs: Information systems Information retrievalccs: Information systems Retrieval models and rankingccs: Information systems Language models
@@ 3.2. Auxiliary Translation Tasks
$3.$ Retrieval-based Query Augmentation
@@ 3.5. Online Relevance Model
Based on the state-of-the-art CSRM-LLM (Stage-2), we distill a twin-tower model for online serving. The model architecture is similar to Que2Search (Liu et al., 2021) with XLM-Roberta-based tokenizer/encoder. The embedding layers are between the query and item towers, while other parameters are tuned separately. In the inference phase, the cosine similarity $c_{q,i}$ of the query and item towers’ outputs will be used as the final prediction score. Note that $c_{q,i}$ ranges from 1-1 to 11, which is not a valid probability applicable in computing KD loss. Following the strategy in  (Liu et al., 2021), we set $\hat{p}_{q,i}=\sigma(s\cdot c_{q,i})$, where $\sigma$ denotes the Sigmoid function, and $s$ is a hyper-parameter set as 88 empirically. We choose the twin-towered structure as it is more friendly for CPU serving. All product embeddings are pre-calculated and refreshed regularly. When receiving an online request, the query embedding is either generated in real time or directly fetched from online cache. For each query, we sample $M+N$ products to construct the query-product pairs for knowledge distillation, including: (1) Top $M$ impressed products from historical logs in recent one month. This data will guide the model to learn the relevance matching of popular query-product pairs. (2) Top NN products retrieved by the EBR model trained by the sparse click and conversion data collected in the new market. EBR candidates are included because the top impressed products are mostly relevant products, while we need more substitute and irrelevant items to balance the training data distribution.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systemsccs: Information systems Information retrievalccs: Information systems Retrieval models and rankingccs: Information systems Language models
@@ 3.2. Auxiliary Translation Tasks
3.3. Retrieval-based Query Augmentation
@@ 3.5. Online Relevance Model
Based on the state-of-the-art CSRM-LLM (Stage-2), we distill a twin-tower model for online serving. The model architecture is similar to Que2Search (Liu et al., 2021) with XLM-Roberta-based tokenizer/encoder. The embedding layers are between the query and item towers, while other parameters are tuned separately. In the inference phase, the cosine similarity $c_{q,i}$ of the query and item towers’ outputs will be used as the final prediction score. Note that $c_{q,i}$ ranges from 1-1 to $1$, which is not a valid probability applicable in computing KD loss. Following the strategy in  (Liu et al., 2021), we set $\hat{p}_{q,i}=\sigma(s\cdot c_{q,i})$, where $\sigma$ denotes the Sigmoid function, and $s$ is a hyper-parameter set as $8$ empirically. We choose the twin-towered structure as it is more friendly for CPU serving. All product embeddings are pre-calculated and refreshed regularly. When receiving an online request, the query embedding is either generated in real time or directly fetched from online cache. For each query, we sample $M+N$ products to construct the query-product pairs for knowledge distillation, including: (1) Top $M$ impressed products from historical logs in recent one month. This data will guide the model to learn the relevance matching of popular query-product pairs. (2) Top $N$ products retrieved by the EBR model trained by the sparse click and conversion data collected in the new market. EBR candidates are included because the top impressed products are mostly relevant products, while we need more substitute and irrelevant items to balance the training data distribution.